OpenStudy (anonymous):

Simplify: (3x^2y^-3)^2 (2x^5y^-2)^-3

OpenStudy (anonymous):

you mean\[(3x ^{2}y ^{-3})^{2} \times (2x ^{5} y ^{-2})^{-3}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[= 3^{2}x ^{4}y ^{-6} \times 2^{-3} x ^{-15}y ^{6}\] \[=(9/8 )\times x ^{-11}\] i think now you can solve this :)

OpenStudy (anonymous):

How do you get (9/8)

OpenStudy (anonymous):

3^2=9 2^-3=1/8 so 9 times 1/8 is 9/8